# Stable Score Distillation for High-Quality 3D Generation

### 14 Dec 2023 v1

@@ 4.3.2 Intrinsic Limitations
Figure 4: Comparisons on 2D optimization. The experiments are run with more training iterations than normally used to expose the intrinsic properties of the estimators. Objects’ local features are more plausible with SSD.
@@
Over-smoothness. $\hat{\boldsymbol{\boldsymbol{\epsilon}}_{\phi}}(\boldsymbol{x}_{t},t,y)=-\sigma_{t}\nabla_{\boldsymbol{x}_{t}}log\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)$ according to Eq. 2. During the 3D optimization process, this term directs $\mathbb{E}_{\epsilon}[\boldsymbol{x}_{t}]=\alpha_{t}\boldsymbol{x}$ towards modes of $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)$. When $t$ is small (e.g. $\leq 200$), this behaviour is precisely what we aim for since $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)\approx\mathbb{P}_{\phi}(\boldsymbol{x};y,t)$ and the mode-seeking term guides $\theta$ such that its rendering $g(\theta,c)$ converges towards the modes of $\mathbb{P}_{\phi}(\boldsymbol{x};y,t)$. However, for large values of $t$, the situation is more complicated. For clarity, let’s consider two modes in $\mathbb{P}_{\phi}(\boldsymbol{x};y,t)$ located at $\boldsymbol{o}_{1}$ and $\boldsymbol{o}_{2}$. At diffusion timestep $t$, they induce modes of $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)$ at $\alpha_{t}\boldsymbol{o}_{1}$ and $\alpha_{t}\boldsymbol{o}_{2}$, respectively, which we term as induced modes. Notably, $\alpha_{t}<1$, resulting in the induced modes being closer to each other than $\boldsymbol{o}_{1}$ and $\boldsymbol{o}_{2}$. As $t$ gradually increases, a new mode in $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y)$, hypothetically located at $\alpha_{t}\boldsymbol{o}_{tr}$, emerges between $\alpha_{t}\boldsymbol{o}_{1}$ and $\alpha_{t}\boldsymbol{o}_{2}$, referred to as a transient mode. Note that this implies $\boldsymbol{o}_{tr}$ lies in between $\boldsymbol{o}_{1}$ and $\boldsymbol{o}_{2}$. With sufficiently large $t$, the transient mode will have higher probability density than the induced ones, causing the mode-seeking term to seek the transient modes. In other words, the 3D rendering $\boldsymbol{x}$ optimizes for $\alpha_{t}\boldsymbol{x}\rightarrow\alpha_{t}\boldsymbol{o}_{tr}$, or equivalently $\boldsymbol{x}\rightarrow\boldsymbol{o}_{tr}$, when $t$ is large. We call $\boldsymbol{o}_{tr}$ a trapping point in $\mathbb{P}_{\phi}(\boldsymbol{x};y)$. Furthermore, as presented in Sec. 3.4, the supervision loss proposed by DreamFusion [30] weights the KL divergences under different $t$ according to $w(t)\frac{\sigma_{t}}{\alpha_{t}}$, which is typically an increasing function with respect to $t$. For example, in DreamFusion, $w(t)=\sigma_{t}^{2}$ and $w(t)\frac{\sigma_{t}}{\alpha_{t}}$ equals to 0 and +$\infty$+ ∞ when $t=0$ and $T$, respectively. Such weighting encourages the mode-seeking term to emphasize on large $t$ and consequently prioritize pursuing the transient modes. In other words, the overall optimization process drives $\boldsymbol{x}=g(\theta,c)$ to approach $\boldsymbol{o}_{tr}$, which lies between the conditional modes of $\mathbb{P}_{\phi}(\boldsymbol{x};y,t)$. Since the timestep $t$ is randomly sampled, the optimization process frequently revisits large $t$ and gets $\boldsymbol{x}$ trapped at $\boldsymbol{o}_{tr}$. This results in over-smoothness as $\boldsymbol{x}$ converges to a middle point of different plausible image modes. One can refer to Fig. 3 (right) for an illustration of the transient-mode problem and the learning trajectory of $\alpha_{t}\boldsymbol{x}$. Note that the transient mode causes $\boldsymbol{x}$ to converge to the middle of two conditional modes as shown in Fig. 3 (left).
It is worth noting that the transient-mode trap underscores a distinction between optimization-based content generation and sampling-based content generation. In the latter case, while it is plausible for $\boldsymbol{x}_{t}$ to reside in a transient mode $\alpha_{t}\boldsymbol{o}_{tr}$, subsequent sampling steps will only decrease $t$. With $t$ decreasing, the transient modes gradually vanishes and the sequence $\boldsymbol{x}_{t-1},\boldsymbol{x}_{t-2},\cdots$ converges to induced modes $\alpha_{\tau}\boldsymbol{o}_{i}$, where $\boldsymbol{o}_{i}$ are conditional modes in $\mathbb{P}_{\phi}(\boldsymbol{x};y)$ and $\tau$ assumes the ordered sequence of $t-1,t-2,...$, progressively approaching the modes of $\mathbb{P}_{\phi}(\boldsymbol{x};y)$. There is no longer persistent force that prioritizes the final generation $\boldsymbol{x}_{0}$ to pursue transient modes, namely $\alpha_{t}\boldsymbol{x}_{0}\approx\alpha_{t}\boldsymbol{o}_{tr}$.
@@ 4.4.1 A Naive Approach
Consider $\boldsymbol{b}=\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)-\boldsymbol{\epsilon}$, same as SDS with $\omega=0$. It may initially seem intuitive that $\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)\approx\boldsymbol{\epsilon}$ due to the denoising nature of training $\hat{\boldsymbol{\epsilon}}_{\phi}$ and thus $\boldsymbol{b}$ induces low variance, which is also an unbiased estimator for $\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)$ as $\mathbb{E}_{\boldsymbol{\epsilon}}[\boldsymbol{\epsilon}]=\boldsymbol{0}$. However, analysis from Sec. 4.3 reveals that the scale and direction of the mode-seeking term are highly dependent on the diffusion timestep $t$ and therefore $\boldsymbol{\epsilon}$ does not always reduce variance as expected. For example, when $t=0$ where $\alpha_{t}=1$ and $\sigma_{t}=0$, we have $\hat{\boldsymbol{\epsilon}}_{\phi}=\boldsymbol{0}$ and the “variance reduction” term in $\boldsymbol{b}$ is actually the only source of variance! We illustrate this observation in Fig. 5. Consequently, it becomes important to reduce the variance of $\hat{\boldsymbol{\epsilon}}_{\phi}$ through a more meticulous design.
Figure 5: Visualization of variance-reduction schemes for gradients. When $t$ is large, e.g. 500, the adaptive variance-reduction term, $c$, is close to 1 and our variance reduction scheme performs similarly to SDS-like variance reduction scheme. On the other hand, when $t$ is small, our method significantly differs from SDS and produces smoother gradients. Note that the SDS-like variance-reduction scheme does not effectively reduce variance of the conditional score for small $t$.
@@ 4.4.2 Adaptive Variance-Reduction
We aim to reduce the variance of $\hat{\boldsymbol{\epsilon}}_{\phi}$ more efficiently by conditioning the variance-reduction term on $\boldsymbol{x}$ and $t$:
$\hat{\boldsymbol{\epsilon}}_{\phi}-c(\boldsymbol{x},t)\boldsymbol{\epsilon},$ (5)
which remains an unbiased estimator of $\hat{\boldsymbol{\epsilon}}_{\phi}$ because the introduced coefficient $c(\boldsymbol{x},t)$ is independent of $\boldsymbol{\epsilon}$, where $c$ is optimally chosen to maximally reduce the variance:
@@
For ease of mathematical analysis, in this work we adopt total variance as our variance definition in Eq. 6 for its wide usage [47]. We define $\Sigma_{XY}$ as the covariance matrix between random vectors $X$ and $Y$, and define $\Sigma_{X}$ as a shorthand for $\Sigma_{XX}$. Then Eq. 6 becomes $c(\boldsymbol{x},t)=\text{argmin}_{k}\ \text{tr}(\Sigma_{\hat{\boldsymbol{\epsilon}}_{\phi}-k\boldsymbol{\epsilon}})$, which has a closed-form solution (see App. Sec. 11 for proof):
@@
Note that $r$ is defined this way to minimize the norm of $\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)-k\boldsymbol{\epsilon}$, namely $r(\boldsymbol{x}_{t},\boldsymbol{\epsilon},y,t)=\text{argmin}_{k}||\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)-k\boldsymbol{\epsilon}||.$ Our experiments reveal that $r$ is quite robust to $\boldsymbol{\epsilon}$ for any combination of $\boldsymbol{x}$, $y$, and $t$, and $c$ is tightly bounded by the extremes of $r$, as shown in App. Fig. 20. Therefore in our experiments we use $r$ as a computationally convenient proxy for $c$. Formally, we define a variance-reduced mode-seeking estimator as:
@@
It is worth noting that unlike Eq. 5, Eq. 9 is no longer an unbiased estimator of $\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)$ since $r$ is now dependent on $\boldsymbol{\epsilon}$. However, in practice, $r$ and $c$ are quite close to each other and the approximation of Eq. 5 with Eq. 9 does not result in any noticeable performance degradation.
@@ 4.5 Augmenting the Mode-Disengaging Term with Low-Variance Mode-Seeking Behaviour
We summarize the 3D generation supervision terms and their behaviour under different regimes of $t$ in Tab. 1. Evidently, we should combine the merits of the mode-disengaging and the mode-seeking terms under different timestep regimes, which leads to our Stable Score Distillation (SSD) estimator. The design is quite simple: when $t>M$, where $M\geq 0$ is a pre-defined timestep threshold, we employ the mode-disengaging term for fast geometry formation. Otherwise, we utilize the variance-reduced mode-seeking term as defined in Eq. 9 to guide the 3D model’s renderings towards plausible image distribution. Notably, we further address the scale mismatch between the two supervision signals as shown in Fig. 22 by scaling the variance-reduced mode-seeking term to match the scale of the mode-disengaging one. The detailed algorithm is provided in Algorithm 1.
@@
large good geometry details over-smoothness
small color-saturation plausibility
@@
3:Input: generation condition, e.g. text prompt $y$
4:Input: CFG scale $\omega$, timestep threshold $M$
5:Input: pre-trained denoiser $\hat{\boldsymbol{\epsilon}}_{\phi}$, e.g. Stable Diffusion
@@
7:$h=\omega\cdot(\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},t,y)-\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},t,\emptyset))$
@@
4.6 Connection with Common Observations and Practices in 3D Content Generation
  • Large CFG Scales. As illustrated in App. Fig. 22, the norms of the mode-disengaging term and the mode-seeking term are vastly different. Consequently, a large CFG scale is necessary to make the scale of the main learning signal $\omega\cdot h$ at least comparable to that of the mode-seeking term. Otherwise the SDS estimator in Eq. 3 would be dominated by the mode-seeking term, which can only generate over-smoothed contents for reasons discussed in Sec. 4.3.2, as shown in App. Fig. 9.
  • Color Saturation. This challenge has been repeatedly observed in text-to-3D generation [30, 21, 10]. Recent work of DreamTime [13] also observes severe color saturation tendency when $t<100$. According to our analysis in Sec. 4.2, when $t$ is small the mode-disengaging term tends to drive the rendering $g(\theta,c)$ away from modes of the natural image distribution, making the rendering implausible. As pointed out in [6], in practice the diffusion model tends to influence fine details and colors, but not large-scale geometry, when $t$ is small. Thus with small $t$ the mode-disengaging term generates implausible colors, without corrupting the general geometry of the objects.
@@ 5.1 Implementation Details
Unless otherwise specified, our experiments are conducted with the threestudio codebase [9]. The high-resolution text-to-NeRF results as shown in Fig. 2 are generated with a batch size of 12 in the low-resolution stage and 1 in the high-resolution stage. All the other 3D assets presented in Fig. 6 and Fig. 7 are generated with batch size 1. We employ Stable Diffusion [34] v2.1-base as our score estimator $\hat{\boldsymbol{\epsilon}_{\phi}}$.
We also evaluate the compatibility between our SSD estimator and other 3D representation, i.e., Gaussian Splatting [17]. We employ the codebase open-sourced by LucidDreamer [19], but disable its DDIM-inversion and replace its interval score matching with SSD. In our experiments, disabling the DDIM inversion speeds up the training process by 75%. The batch size is set to 4 as used in LucidDreamer. Since the Perp-Neg [1] implemented in the codebase is able to alleviate the issue of color saturation to some extent, where the negative prompt is “unrealistic, blurry, low quality, out of focus, ugly, low contrast, dull, dark, low-resolution, oversaturation”, we set $M=100$. The Gaussian Splatting generations are presented in Fig. 1.
5.2 Illustrative Example
First, we evaluate the learning behaviour of the mode-seeking, mode-disengaging and SSD terms with an illustrative example, as shown in Fig. 3. In this experiment, we assume that $\boldsymbol{x}=\theta\in\mathbb{R}^{2}$, and $\mathbb{P}_{\phi}(\boldsymbol{x})=0.2\mathcal{N}([0,0]^{T},0.1I)+0.4\mathcal{N}([1,1]^{T},0.05I)+0.4\mathcal{N}([2,1]^{T},0.05I)$. The two Gaussian components located at $[1,1]^{T}$ and $[2,1]^{T}$ are conditional modes, while the one at $[0,0]^{T}$ is a singular one (refer to Sec. 4.1 for definitions). We initialize $\theta$ at $[0,0]^{T}$ to simulate the common situation where the 3D generation algorithms start with empty renderings. We then substitute the mode-seeking, mode-disengaging and our SSD estimators for the SDS estimator in Eq. 3 one by one, and record the learning trajectory of $\theta$. To validate the trap-escaping property of the mode-disengaging term we initialize a $\theta$ to the trapping point (see Sec. 4.3.2), namely $[1.5,1]^{T}$, and use $h$ to supervise $\theta$ afterwards. The learning trajectories are illustrated in Fig. 3 (left). We observe that the mode-seeking trajectory gets trapped between the two conditional modes rapidly, not converging to any specific conditional mode of $\mathbb{P}_{\phi}(\boldsymbol{x})$. To inspect whether the issue is caused by the transient-mode problem, we present the density map of $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y)$ at $t=350$, and the learning trajectory of $\alpha_{t}\boldsymbol{x}$ in Fig. 3 (right). The density map reveals that the two conditional modes of $\mathbb{P}_{\phi}(\boldsymbol{x};y)$ result in a transient mode in $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y)$, whose probability density is higher than the induced modes. Consequently $\alpha_{t}\boldsymbol{x}_{t}$ indeed approaches the transient mode $\alpha_{t}\boldsymbol{o}_{tr}$. Conversely, the mode-disengaging term can propel $\theta$ away from the trapping point $\boldsymbol{o}_{tr}$. However, although at the beginning it guides $\theta$ towards a conditional mode, it ultimately steers $\theta$ into a low-density region of $\mathbb{P}_{\phi}(\boldsymbol{x})$. In contrast, the proposed SSD swiftly guides $\theta$ towards a conditional mode, avoids getting trapped, and finally converges to a point with high density in $\mathbb{P}_{\phi}(\boldsymbol{x})$.

### 07 Feb 2024 v2

@@ 4.3.2 Intrinsic Limitations
Over-smoothness. According to proof (A.4) in DreamFusion [24], the mode-seeking term drives $\mathbb{P}(\boldsymbol{x}_{t}|\boldsymbol{x};y,t)=\mathcal{N}(\alpha_{t}\boldsymbol{x},\sigma_{t}^{2}I)$ to align with the high-density region of $p_{\phi}(\boldsymbol{x}_{t};y,t)$. Naturally, this means that during the 3D optimization process, this term directs the mean of the above Gaussian distribution, namely $\alpha_{t}\boldsymbol{x}$, towards modes of $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)$. When $t$ is small (e.g., $\leq 200$), this behaviour is precisely what we aim for since $\mathbb{P}_{\phi}(\boldsymbol{x}_{t};y,t)\approx\mathbb{P}_{\phi}(\boldsymbol{x};y,t)$ and the mode-seeking term guides $\theta$ such that its rendering $g(\theta,c)$ converges towards the modes of $\mathbb{P}_{\phi}(\boldsymbol{x};y,t)$. However, for large values of $t$, the situation is more complicated. For clarity, let’s consider two modes in $\mathbb{P}_{\phi}(\boldsymbol{x};y,t)$ located at $\boldsymbol{o}_{1}$ and $\boldsymbol{o}_{2}$. Ideally, we would like $\boldsymbol{x}$ to converge around them. Equivalently, we would like $\alpha_{t}\boldsymbol{x}\approx\alpha_{t}\boldsymbol{o}_{1}$ or $\alpha_{t}\boldsymbol{x}\approx\alpha_{t}\boldsymbol{o}_{2}$ for any $t$. We call $\alpha_{t}\boldsymbol{o}_{*}$ as induced modes. However, when $t$ gets high, $\alpha_{t}$ is rather low. And now the induced modes gets so close that they ”melt” into a single mode which lies between the induced ones. We assume such a mode is $\alpha_{t}\boldsymbol{o}_{tr}$, which we call transient mode. Trivially, this means $\boldsymbol{o}_{tr}$ lies between $\boldsymbol{o}_{1}$ and $\boldsymbol{o}_{2}$. Thus the mode-seeking term drives $\alpha_{t}\boldsymbol{x}$ to seek the transient modes $\alpha_{t}\boldsymbol{o}_{tr}$, instead of the induced ones as expected, when $t$ is large. In other words, the 3D rendering $\boldsymbol{x}$ optimizes for $\boldsymbol{x}\rightarrow\boldsymbol{o}_{tr}$. We call $\boldsymbol{o}_{tr}$ a trapping point in $\mathbb{P}_{\phi}(\boldsymbol{x};y)$. Since the timestep $t$ is randomly sampled during optimization, the overall optimization process frequently revisits large $t$ and gets $\boldsymbol{x}$ trapped at $\boldsymbol{o}_{tr}$. This results in over-smoothness as $\boldsymbol{x}$ now converges to a middle point of plausible image modes, trying to “average” the plausible contents. See Fig. 7 for the visualization of the mode-seeking term causing model parameters to be trapped by transient modes.
@@
Connection with common observations and practices in 3D content generation. We find that many observations and practices of various frameworks have non-trivial connections with the previous analysis. Interested ones can refer to Sec. 12 for details.
@@ 4.4.1 A Naive Approach
Consider $\boldsymbol{b}=\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)-\boldsymbol{\epsilon}$, same as SDS with $\omega=0$. It may initially seem intuitive that $\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)\approx\boldsymbol{\epsilon}$ due to the denoising nature of training $\hat{\boldsymbol{\epsilon}}_{\phi}$ and thus $\boldsymbol{b}$ induces low variance, which is also an unbiased estimator for $\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)$ as $\mathbb{E}_{\boldsymbol{\epsilon}}[\boldsymbol{\epsilon}]=\boldsymbol{0}$. However, analysis from Sec. 4.3 reveals that the scale and direction of the mode-seeking term are highly dependent on the diffusion timestep $t$ and thus $\boldsymbol{\epsilon}$ does not reduce variance always. For example, when $t=0$ where $\alpha_{t}=1$ and $\sigma_{t}=0$, we have $\hat{\boldsymbol{\epsilon}}_{\phi}=\boldsymbol{0}$ and the “variance reduction” term in $\boldsymbol{b}$ is actually the only source of variance! We illustrate this observation in Fig. 9. Consequently, it becomes important to reduce the variance of $\hat{\boldsymbol{\epsilon}}_{\phi}$ through a more meticulous design.
@@ 4.4.2 Adaptive Variance-Reduction
We aim to reduce the variance of $\hat{\boldsymbol{\epsilon}}_{\phi}$ more efficiently by conditioning the variance-reduction term on $\boldsymbol{x}$ and $t$: $\hat{\boldsymbol{\epsilon}}_{\phi}-c(\boldsymbol{x},t)\boldsymbol{\epsilon}$, which remains an unbiased estimator of $\hat{\boldsymbol{\epsilon}}_{\phi}$ because the introduced coefficient $c(\boldsymbol{x},t)$ is independent of $\boldsymbol{\epsilon}$, where $c$ is optimally chosen to maximally reduce the variance:
@@
Based on some mathematical treatments as detailed in Sec. 8, we instead optimize the objective function with $r$, which is a computationally cheap proxy for the optimal $c$ and defined as:
@@
Formally, we define a variance-reduced mode-seeking estimator as:
@@
It is worth noting that Eq. 7 is no longer an unbiased estimator of $\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},y,t)$ since $r$ is now dependent on $\boldsymbol{\epsilon}$. However, in practice, $r$ and $c$ are quite close to each other and the approximation with Eq. 7 does not result in any noticeable performance degradation. See Fig. 18 for numerical confirmation.
@@ 4.5 Augmenting the Mode-Disengaging Term with Low-Variance Mode-Seeking Behaviour
We summarize the 3D generation supervision terms and their behaviour under different regimes of $t$ in Tab. 1. Evidently we should combine the merits of the mode-disengaging and the mode-seeking terms under different timestep regimes, which leads to our Stable Score Distillation (SSD) estimator. The design is quite simple: when $t>M$, where $M\geq 0$ is a pre-defined timestep threshold, we employ the mode-disengaging term for fast content formation. Otherwise we utilize the variance-reduced mode-seeking term as defined in Eq. 7 to guide the 3D model’s renderings towards plausible image distribution. Notably, we further address the scale mismatch between the two cases by scaling the variance-reduced mode-seeking term to match the scale of the mode-disengaging one, inspired by the scale mismatch between the terms as shown in Fig. 19. The detailed algorithm is provided in Algorithm 1.
@@
large trap escaping over-smoothness
small implausibility (e.g., floaters) plausibility
@@
3:Input: generation condition, e.g., text prompt $y$
4:Input: Timestep threshold $M$
5:Input: pre-trained denoiser $\hat{\boldsymbol{\epsilon}}_{\phi}$, e.g., Stable Diffusion
@@
7:$h=\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},t,y)-\hat{\boldsymbol{\epsilon}}_{\phi}(\boldsymbol{x}_{t},t,\emptyset)$
@@ 5.1 Implementation Details
Our experiments are mainly conducted with the threestudio codebase [8]. To save vram usage, during the first 5000 steps the training resolution is 64. The batch size for low-resolution and high-resolution stages are 8 and 1 respectively. We employ Stable Diffusion [28] v2.1-base as our score estimator $\hat{\boldsymbol{\epsilon}_{\phi}}$. We employ the ProgressBandHashGrid with 16 levels implemented in the codebase as our 3D representation. The learning rates are 0.01 for the encoding and 0.001 for the geometry networks. We conduct our experiments with one A100 80GB on Ubuntu 22.04LTS. Each asset is trained for 15000 steps. Also, during the first 5000 steps we sample $t$ from [20, 980] and anneal it to [20, 500] afterwards. The loss is SSD loss, plus the orient loss in the codebase with a coefficient 1000. Note that our method does not need to tune CFG scales like previous works [24].
5.2 Evaluation on Numerical Experiment
First, we demonstrate the efficacy of our SSD as a general-purpose estimator for mode approximation with a simple numerical experiment on a mixture-of-Gaussian distribution. Note that the mixture-of-Gaussian distribution is a general-purpose distribution approximator in that it can approximate any continuous distribution with any precision [7]. Thus it suffices to evaluate our SSD on mixture-of-Gaussian. See Sec. 13 for details.